Two years ago a woman was 7 times as old as her daughter but in 3 years time she wound be only 4 times as old as the girl .how old are they now?
Found 2 solutions by josgarithmetic, ikleyn:
Answer by josgarithmetic(39616)
Answer by ikleyn(52776)

.

    x = daughter 

    7x = mother


In three years

    daughter = (x+3)

    mother =   7x + 3


Equation

    7x + 3 = 4*(x+3).


Solution

    7x + 3  = 4x + 12

    7x - 4x = 12 -3

    3x      = 9

     x      = 9/3 = 3.


ANSWER.  Daughter is 3 years old.   Mother is  7*3 = 21 years old.

Solved.

--------------------

There is a bunch of lessons on age word problems
    - Age problems and their solutions
    - A fresh formulation of a traditional age problem
    - Really intricate age word problems
    - Selected age word problems from the archive
    - Age problems for mental solution
    - Miscellaneous age problems
in this site.

Read them and become an expert in solving age problems.

Also,  you have this free of charge online textbook in ALGEBRA-I in this site
    - ALGEBRA-I - YOUR ONLINE TEXTBOOK.

The referred lessons are the part of this online textbook under the topic "Age word problems".


Save the link to this online textbook together with its description

Free of charge online textbook in ALGEBRA-I
https://www.algebra.com/algebra/homework/quadratic/lessons/ALGEBRA-I-YOUR-ONLINE-TEXTBOOK.lesson

to your archive and use it when it is needed.